OpenStudy (anonymous):

Which of the following is a correct equation for the line passing through the point (-2,1) and having slope m = 1/2?

OpenStudy (campbell_st):

the slope intercept equation is y = mx + b = m gradient and b is the y-intercept so in your question you have y = 1/2 x + b substitute x = -2 and y = 1 into the equation above to find b then you'll have the equation of your line.